What is the energy change when 21.0 grams of water cools from 53.0 °C to 20.5 °C?
Answer #1

Answer

By using the equation Q= m*c*deltaT

Q= energy = ?     ;  m = mass of water= 21 grams   

c= specific heat of water= 4.184 J/gC

deltaT= temperature change = 53-20.5= 32.5 degree C

now

Q= 21*4.184*32.5=2855.6 J